Grade 9Mechanics

two balls are dropped from same height at 1 second time interval.The seperation between the two balls after 3 seconds is

Mayank Saxena
9 Years agoGrade 9

1 Answer

Vikas TU
9 Years ago
Dear Student,
As the another is discharged from a similar stature a moment after the first. That implies first body travel 2 sec while the second body travel just for 1 sec. 
For the primary body, 
S = ut + ½ at^2 
S = 0 + ½ *g* 4 
S = 2g m 
For the second body, 
S= ut+1/2 at^2 
S =1/2 g 
Contrast between the separation = 3/2g = 15m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)